OpenStudy (anonymous):

the measure of angle t is 35 degrees greater than 4 times the measurement of its supplement what is the supplement of angle t

OpenStudy (anonymous):

if you call angle supplement x then: x+t=180 (by definition of supplemental angle) also you know that t=35+4x so you have system of equations: x+t=180 t=35+4x can you solve it?

OpenStudy (anonymous):

substitute value of t (from the second equation) to the first: x+35+4x=180 5x=145 x=29 (this is the value of supplemental angle) to find the value of t: t=180-29 t=151 (degrees)